find three consecutive positive numbers such that the product of the first and third, minus the second,is 1 more than 3 times the third
the smallest integer is __ , the second smallest integer is ___, the largest integer is ___



thanks

Answer by HasanSahin(52)

one of the classics..observe..
Let's say x,x+1,x+2 for the consecutive integers..
x*(x+2)-(x+1) = 3*(x+2)+1
x^2+2x-x-1 = 3x+6+1
x^2-2x-8 = 0
(x-4)*(x+2) = 0
x={4,-2}
You must choose as x = 4 because the integers are positive..
So that,4,5,6 are the integers..
RF.